OpenStudy (anonymous):

Keenan has 28 CDs and 44 DVDs. What is the ratio of DVDs to CDs? 11:7 28:1 7:11 1:28 @hartnn @paki @Hero @op13456

hartnn (hartnn):

DVD/ CD = 44/28 reduce this! both are divisible ny 4

hartnn (hartnn):

**by 4

OpenStudy (anonymous):

so its c

hartnn (hartnn):

nopes

OpenStudy (anonymous):

a?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

Check the order yes a

hartnn (hartnn):

44 = 4 *11 28 = 4*7

OpenStudy (anonymous):

In questions like these you must check the order in which they are asking Like they asked DVDs to CDs

OpenStudy (anonymous):

DVDS to CDS 44:28 just simplify down and it's 11:7 (hoped this helped)

OpenStudy (anonymous):

thx